# Scaling Analysis of Interleaved Speech-Text Language Models

### v1

@@ A.2 Hyper-Parameters
We train all interleaved SLMs with the same settings - context length of 2048, AdamW optimiser, with a Cosine decay learning rate scheduler with a warmup of 1% of training steps and a final minimal learning rate of 5e-5. We use bfloat16 precision, with flash-attention2 and data packing. We preform gradient normalisation with a norm of 0.5. We balance the modalities - text-only, speech-only and interleaved by mixing samples based on length distribution of each dataset, and not on exact tokens per batch. We use a maximum learning rate of 5e-4 for most models, but lower this to 3e-4 for training of compute as the longer step count meant that the learning rate “shoulder” was high for too long and induced instability in training. For the compute budget we considered both 3e-4 and 5e-4 for all 4 models, and used the best for each by validation loss. Specifically this was 3e-4 for Qwen2.5-7B, and 5e-4 for all others (albeit not by a big difference).
@@ A.3 Evaluation
For generative perplexity - we use the first 1000 samples of s-StoryCloze (correct samples only), and take the first 3 seconds as a prompt. We then generate a maximum of 150 tokens with top-K=25, and a temperature of 0.8, which matches Slam (Maimon et al., 2025a). We transcribe the samples using Whisper-large-v-turbo model (Radford et al., 2023) and measure Perplexity using the Llama--B model (Grattafiori et al., 2024).
@@ Appendix B Additional Results
We provide here additional results in analysing SLM performance. Specifically, in Figure 5 we show the impact of interleaving and TWIST initialisation for OPT125, which is a smaller model compared to Qwen2.5-0.5B used in Figure 3. We note that for this model interleaving under-performs compared to speech-only for all but the largest budget (where they are comparable). We also note that the non-TWIST initialised model (GSLM) closes the gap quite substantially, unlike with Qwen.

### v2

@@ A.2 Hyper-Parameters
We train all interleaved SLMs with the same settings - context length of , AdamW optimiser, with a Cosine decay learning rate scheduler with a warm-up of of training steps and a final minimal learning rate of . We use bfloat16 precision, with flash-attention2 and data packing. We preform gradient normalisation with a norm of . We balance the modalities - text-only, speech-only and interleaved by mixing samples based on length distribution of each dataset, and not on exact tokens per batch. We use a maximum learning rate of for most models, but lower this to for training of compute as the longer step count meant that the learning rate “shoulder” was high for too long and induced instability in training. For the compute budget we considered both and for all 4 models, and used the best for each by validation loss. Specifically this was for Qwen-B, and for all others (albeit not by a big difference).
@@ A.3 Evaluation
For generative perplexity - we use the first 1000 samples of sSC (correct samples only), and take the first 3 seconds as a prompt. We then generate a maximum of tokens with , and a temperature of , which matches Slam (Maimon et al., 2025a). We transcribe the samples using Whisper-large-v-turbo model (Radford et al., 2023) and measure Perplexity using the Llama--B model (Grattafiori et al., 2024).
@@ Appendix B Additional Results
B.1 When Does Text Benefit Speech Modelling?
We provide here additional results in analysing SLM performance. Specifically, in Figure 5 we show the impact of interleaving and TWIST initialisation for OPT, which is a smaller model compared to Qwen-B used in Figure 3. We note that for this model interleaving under-performs compared to speech-only for all but the largest budget (where they are comparable). We also note that the non-TWIST initialised model (GSLM) closes the gap quite substantially, unlike with Qwen. This could further the support the need to use strong TextLM model families to achieve good scaling performance.
B.2 Diverse Speaker Analysis
In order to analyse how much results vary per-speaker in our suggested multi-speaker benchmark, we report per-speaker metrics for select models. We provide the results in Table 5. Results suggest a slight harm to performance for the British accent, which demonstrates the important of evaluating with diverse speakers. However, the difference in performance across speakers is limited, which shows the robustness of our models.
For evaluation efficiency, we only selected four speakers for the official multi-speaker benchmark. However, in order to further establish to what extent these four speakers represent expected performance on other speakers we select four more speakers. We provide results in Table 6, which show similar trends and scores to the original speakers.
| sSC/tSC | af_bella | am_puck | bf_emma | bm_george |
| SIMS (Q1.5) | 56.23/84.45 | 57.35/83.59 | 54.30/82.68 | 55.32/84.77 |
| SIMS (L3) | 59.27/86.48 | 59.49/86.85 | 57.19/85.03 | 56.97/86.53 |
| SIMS (Q7) | 61.41/88.35 | 62.37/88.62 | 58.52/87.28 | 59.01/88.78 |
| sSC/tSC | af_heart | am_michael | bf_isabella | bm_fabel |
| SIMS (Q1.5) | 56.81/85.46 | 56.76/85.03 | 55.64/81.24 | 55.53/81.24 |
| SIMS (L3) | 58.90/87.23 | 58.15/87.65 | 56.97/83.86 | 58.20/84.61 |
| SIMS (Q7) | 61.52/90.22 | 61.57/90.06 | 58.52/85.94 | 59.27/85.57 |
B.3 Scaling Analysis